# Transforming hyperspectral images into chemical maps: A new deep learning based approach to hyperspectral image processing

### v1

@@ 3.1 U-Net Convergence
The graphs’ noisy nature was attributed to the relatively small dataset size compared to the rather large U-Net. Additionally, a batch size of one was used due to memory constraints, which may also lead to noisy gradients.
@@ 3.2 Pork belly mean fat predictions
The mean fat content was the primary quality parameter for the predictions. The reference values were given for whole pork bellies, not individual slices. Thus, to get a mean pork belly prediction for U-Net, the pixel-wise predictions for all pixels inside the masks of all slices belonging to a given pork belly were averaged. Conversely, the PLS model was trained on mean belly spectra, and thus, its predictions were already on the complete belly level, and no averaging was required. Alternatively, it was possible with PLS to take an approach similar to the U-Net by averaging the pixel-wise predictions of the chemical maps. This averaging yielded a different result for the PLS predictions. In Fig. 7 shows the belly-wise predictions for U-Net, PLS on mean belly spectra, and PLS averaged over the chemical, and Table 2 summarizes their test set RMSEs.
U-Net outperformed PLS on fat classes F2, F4, and F5, while PLS on mean belly spectra showed the best performance on F1, and PLS averaged over chemical maps showed the best performance on F3. Notice also that the U-Net generally had a slightly lower RMSE on the test set than either PLS method. U-Net showed a minimal discrepancy between RMSE on the CV and test sets compared to PLS. In addition to showing RMSE, the line of best fit was computed for each method for both the CV and the test sets. For these lines, sYX was reported, which is the RMSE around the line of best fit but with two fewer degrees of freedom due to the slope and intercept computations. The sYX values are close to the corresponding RMSE values for both U-Net and PLS on mean belly spectra, indicating well-calibrated models. This calibration measure degenerates slightly for PLS when using the average pixel-wise predictions.
Figure 6: Pork belly-wise mean fat predictions. For U-Net and pixel-wise PLS, the mean fat prediction for each pork belly was computed as the mean over all pixels of all chemical maps belonging to that pork belly. For PLS on mean spectra, the belly-wise mean prediction was obtained directly. (a) U-Net. (b) U-Net with fat classes highlighted. (c) PLS on mean spectra. (d) PLS on mean spectra with fat classes highlighted.
Figure 7: Continued pork belly-wise mean fat predictions. (e) Pixel-wise PLS. (f) Pixel-wise PLS with fat classes highlighted.
@@
PLS on mean spectra 4.75 5.97 6.29 7.06 5.03 5.80
Pixel-wise PLS 5.07 6.06 5.41 6.27 8.08 6.06
Table 2: RMSE (in %) of pork belly-wise fat predictions on the test set for U-Net, PLS on pork belly-wise mean spectra, and pixel-wise PLS.
@@ 3.3 Chemical maps
Interestingly, PLS makes predictions that most resemble spatially uncorrelated noise and massive deviations from the $0-100\%$ range. However, when averaging the otherwise noisy predictions, PLS achieves a mean fat prediction not too far from the reference value as seen in Fig. 7(a). Every PLS chemical map prediction in the dataset, CV and test sets alike, exhibits this phenomenon of noisy chemical maps with somewhat accurate mean values. Conversely, U-Net generated chemical maps that were spatially structured while constraining itself to the $0-100\%$ range and simultaneously achieving a mean prediction close to the reference value. Furthermore, while PLS always makes predictions for which the histogram resembles a normal distribution (such as shown in Fig. 8(e)), U-Net predicts a different distribution for each pork belly slice.
@@ 3.3.1 Spatial distribution analysis
Here, the spatial distribution of pixel-wise predictions is analyzed. First, a qualitative analysis is presented based on the correlation between fat content and firmness of the bellies, measured by the finger pressure method (Maw et al. 35. Then, a statistical analysis of the prediction distribution was conducted.
@@
Fig. 10 plots chemical map predictions on the test set, averaged over each of the 15 regions in the pork bellies, against finger pressure for the same regions. Here, a clear negative correlation exists between predicted fat content and finger pressure. Additionally, U-Net and PLS predict higher fat content for the dorsal cranial sections than for the ventral caudal sections.
Figure 10: Relationship between mean predicted fat percentages, (a) U-Net. (b) PLS, and mean finger pressure for each of the 15 sections of pork bellies in the test set (see regions in Fig. 1). The Pearson correlation coefficients indicate a high negative linear correlation for both models.
Inspired by [26], an analysis of the variance of the spatial distribution of the predicted chemical maps was performed. For this, a predicted chemical map, $\widehat{\mathbf{Y}}$, was considered with its mask, $\mathbf{M}$, and its mean prediction, $\hat{\mathbf{y}}$ (here, the batch index is omitted for readability). The variance ($\sigma^{2}$) of the part of $\widehat{\mathbf{Y}}_{h,w}$ lying inside $\mathbf{M}$ was analyzed. Quantifying how much of the variance was due to spatial correlation was of particular interest. To this end, the so-called nugget effect ($C_{0}$) was computed, which is the semi-variogram ($\gamma^{2}$) evaluated as the pair-wise distance goes to zero (Matheron 34).
@@
Under mild assumptions, subtracting $C_{0}$ from the total variance, $\sigma^{2}$, of $\widehat{\mathbf{Y}}$ inside $\mathbf{M}$ gives the covariance between neighboring pixels (Barnes 11). Thus, $C_{0}$ is a measure of spatially uncorrelated variance. Consequently, $\frac{C_{0}}{\sigma^{2}}$ and $1-\frac{C_{0}}{\sigma^{2}}$ are the ratios of spatially uncorrelated, respectively, correlated variance. Table 3 shows the means of these measures computed over the test set for both PLS and U-Net. Almost all of the variance in U-Net-generated chemical maps was spatially correlated. In contrast, only around three percent of the variance in PLS-generated chemical maps was spatially correlated. The low $C_{0}$ value for U-Net predictions can likely be attributed to the following equivalence between $C_{0}$ 9 and U-Net’s smoothness loss (SL) term Eq. (5): $C_{0}(\widehat{\mathbf{Y}},\mathbf{\mathbf{M}})=\frac{1}{4}SL\left(\widehat{\mathbf{Y}},\mathbf{M}\right)$. As such, U-Net was directly optimized to minimize $C_{0}$, which was not the case for PLS.
@@ 4.2 Bias and scale correction
Consider the U-Net and PLS predictions in Fig. 7. For PLS on mean spectra, the line of best fit for the CV set was guaranteed to be the identity due to the least squares nature of PLS. However, this was not the case for U-Net. Deep learning regressors do not guarantee to find a least squares solution (Igel and Oehmcke 28), especially when training with bulk references and making sample-wise predictions (Engstrøm et al. 18). However, the U-Net’s CV set sYX was very close to its RMSE. Furthermore, bias and scale correction do not change the visual appearance of the chemical maps, nor do they change the values of the ratios in Table 3. The latter points also apply to pixel-wise PLS. For these reasons, no bias or scale corrections were performed.
@@ 4.3 Chemical maps
PLS and the modified U-Net had approximately the same performance when evaluated on the mean fat predictions for pork bellies (Fig. 7). However, as evident by the spatial statistics in Table 3, U-Net generated much more meaningful chemical maps than PLS. Additionally, using the PLS model to predict chemical maps gave mean chemical map predictions that were different from those achieved by the same PLS model evaluated on a mean spectrum as evident by the differences between Fig. 6(c) and Fig. 7(a). U-Net, on the other hand, directly generates a chemical map, and, as such, the mean fat prediction of a belly was given by the mean of the pixel-wise predictions on the slices belonging to that belly. In addition to generating spatially aware, locally smooth chemical maps, U-Net almost always stayed inside the $0-100\%$ range. This was unlike PLS, where most predictions lay outside this range.
@@ 4.4 U-Net extensions
While the goal of this study was to serve as a proof of concept, it was likely possible to further decrease the mean squared error of the U-Net by performing more elaborate studies regarding the values of hyperparameters, including the previously mentioned loss-multipliers, learning rate, optimizer, and all other choices that go into training and evaluating a deep learning model. While such performance optimization was not of interest in this study, it is possible to search for more optimal hyperparameters using, e.g., a Tree-Structured Parzen Estimator (TPE) (Bergstra et al. 13) such as implemented by [1].
In this study, the fat content of pork bellies was analyzed. The dataset includes references for additional parameters (Albano-Gaglio et al. 2). Future developments of this work involve generating chemical maps for multiple parameters using a single model, e.g., a U-Net. A natural extension of the presented loss function would be suitable in this case. First, each parameter should be scaled by its standard deviation to ensure equal contribution to the total loss. Afterward, it is simple to compute an MSE term, Eq. (3), and an SL term, Eq. (5), for each parameter. Interestingly, optimizing the model to regress multiple parameters yields additional possibilities related to the out-of-bounds loss, Eq. (4), that aims to constrain the pixel-wise prediction of each parameter to values between 0 and 100 percent. In addition to constraining each parameter, the sum of parameters should also be constrained in this range. If all parameters can be measured, they should sum to 100 percent. Thus, it is hypothesized that the generalization of the presented work to multiple parameters is not only feasible but perhaps even beneficial compared to the single-parameter case.
Another possible direction for extending this work can be to maintain U-Net’s chemical map generation capability while also having it learn to perform segmentation between foreground (pork belly) and background, a task for which it was initially designed (Ronneberger et al. 39).
@@ 5 Conclusion
This study has proposed a novel end-to-end deep learning approach based on U-Net for chemical map generation using hyperspectral images of pork bellies with associated mean fat reference values as a case study. By training a modified U-Net with a robust optimization scheme and a multi-faceted customized loss function, using only a mean reference value, the model has learned to generate chemical maps with a high degree of spatial correlation and to make predictions strictly inside the 0-100% range of fat values. These findings contrast the current PLS-based pixel-wise approach to chemical map generation that will generate chemical maps with a low degree of spatial correlation and predictions ranging from -1500-1500%, significantly exceeding the realm of the physically possible. Simultaneously, the U-Net-based approach achieves a lower RMSE than PLS when averaging the chemical maps for comparison with the mean reference values. Thus, this study indicates that the U-Net-based approach enables the generation of chemical maps with meaningful spatial features, emphasizing that the analysis of hyperspectral images can benefit from joint consideration of the spectral and spatial features.

### v5

@@ 3.1 U-Net Convergence
The graphs’ noisy nature was attributed to the relatively small dataset size compared to the rather large U-Net. Additionally, like for Ronneberger et al. [41], a batch size of one was used due to memory constraints, which may also lead to noisy gradients. The weight restoration and learning rate reduction handle this instability. In particular, if the validation loss does not improve, the weights are restored to a known good state, and the optimization continues from that point with a reduced learning rate to minimize fluctuations in the loss. Indeed, Fig. 5 shows that the magnitude of fluctuations in the loss decreases as the number of epochs increases, particularly when the learning rate is reduced. The same holds for the other CV splits as shown in  B.
The code for training and evaluating U-Net is released [^4] https://github.com/sm00thix/PorkBellyHSI under the LPGL 3.0 License. The U-Net ensemble’s weights are released [^5] https://huggingface.co/Sm00thix/unet_chemical_map under CC BY-NC 4.0.
@@ 3.2 Pork belly mean fat predictions
The mean fat content was the primary quality parameter for the predictions. The reference values were given for whole pork bellies, not individual slices. Thus, to get a mean pork belly prediction for U-Net, the pixel-wise predictions for all pixels inside the masks of all slices belonging to a given pork belly were averaged. Conversely, the PLS model could be evaluated directly on the pork belly-wise mean spectra. Fig. 6 shows the belly-wise predictions for U-Net and PLS, and Table 2 summarizes their test set RMSEs.  C shows additional metrics.
U-Net outperformed PLS on fat classes F2, F4, and F5, while PLS showed the best performance on F1 and F3. Notice also that the U-Net generally had a slightly lower RMSE on the test set than PLS. U-Net showed a minimal discrepancy between RMSE on the CV and test sets compared to PLS. In addition to showing RMSE, the line of best fit was computed for each method for both the CV and the test sets. For these lines, sYX was reported, which is the RMSE around the line of best fit but with two fewer degrees of freedom due to the slope and intercept computations. The sYX values are close to the corresponding RMSE values for both U-Net and PLS, indicating well-calibrated models.
Figure 6: Pork belly-wise mean fat predictions. (a) U-Net. (b) U-Net with fat classes highlighted. (c) PLS. (d) PLS with fat classes highlighted.
@@
PLS 4.84 5.54 6.12 6.59 5.51 5.64
Reference std. 5.88 6.35 8.38 6.39 6.08 13.2
Reference mean 25.0 29.8 38.4 44.1 60.3 36.1
Table 2: RMSE (in %) of pork belly-wise fat predictions on the test set for U-Net and PLS. The reference means and standard deviations are shown for interpretation of the models’ RMSE values.
@@ 3.3 Chemical maps
Interestingly, PLS makes predictions that most resemble spatially uncorrelated noise and massive deviations from the $0-100\%$ range. However, when averaging the otherwise noisy predictions, PLS achieves a mean fat prediction not too far from the reference value as seen in Fig. 6(c). Every PLS chemical map prediction in the dataset, CV and test sets alike, exhibits this phenomenon of noisy chemical maps with somewhat accurate mean values. Conversely, U-Net generated chemical maps that were spatially structured while constraining itself to the $0-100\%$ range and simultaneously achieving a mean prediction close to the reference value. Furthermore, while PLS always makes predictions for which the histogram resembles a normal distribution (such as shown in Fig. 7(e)), U-Net predicts a different distribution for each pork belly slice.
@@ 3.3.1 Spatial distribution analysis
Here, the spatial distribution of pixel-wise predictions is analyzed. First, a qualitative analysis is presented based on the correlation between fat content and firmness of the bellies, measured by the finger pressure method (Maw et al. 36. Then, a statistical analysis of the prediction distribution was conducted.
@@
Fig. 9 plots chemical map predictions on the test set, averaged over each of the 15 regions in the pork bellies, against finger pressure for the same regions. Here, a clear negative correlation exists between predicted fat content and finger pressure for both U-Net (-0.91) and PLS (-0.96). Additionally, U-Net and PLS predict higher fat content for the dorsal cranial sections than for the ventral caudal sections.
Figure 9: Relationship between mean predicted fat percentages, (a) U-Net, (b) PLS, and mean finger pressure for each of the 15 sections of pork bellies in the test set (see regions in Fig. 1). The Pearson correlation coefficients indicate a high negative linear correlation for both models.
Inspired by [27], an analysis of the variance of the spatial distribution of the predicted chemical maps was performed. For this, a predicted chemical map, $\widehat{\mathbf{Y}}$, was considered with its mask, $\mathbf{M}$, and its mean prediction, $\hat{\mathbf{y}}$ (here, the batch index is omitted for readability). The variance ($\sigma^{2}$) of the part of $\widehat{\mathbf{Y}}_{h,w}$ lying inside $\mathbf{M}$ was analyzed. Quantifying how much of the variance was due to spatial correlation was of particular interest. To this end, the so-called nugget effect ($C_{0}$) was computed, which is the semi-variogram ($\gamma^{2}$) evaluated as the pair-wise distance goes to zero (Matheron 35).
@@
Under mild assumptions, subtracting $C_{0}$ from the total variance, $\sigma^{2}$, of $\widehat{\mathbf{Y}}$ inside $\mathbf{M}$ gives the covariance between neighboring pixels (Barnes 11). Thus, $C_{0}$ is a measure of spatially uncorrelated variance. Consequently, $\frac{C_{0}}{\sigma^{2}}$ and $1-\frac{C_{0}}{\sigma^{2}}$ are the ratios of spatially uncorrelated, respectively, correlated variance. Table 3 shows the means of these measures computed over the test set for both PLS and U-Net. Almost all of the variance in U-Net-generated chemical maps was spatially correlated. In contrast, only around two percent of the variance in PLS-generated chemical maps was spatially correlated. The low $C_{0}$ value for U-Net predictions can likely be attributed to the following equivalence between $C_{0}$, Eq. (9), and U-Net’s smoothness loss (SL) term, Eq. (5): $C_{0}(\widehat{\mathbf{Y}},\mathbf{\mathbf{M}})=\frac{1}{4}SL\left(\widehat{\mathbf{Y}},\mathbf{M}\right)$. As such, U-Net was directly optimized to minimize $C_{0}$, which was not the case for PLS.
@@ 4.2 Bias and scale correction
Consider the U-Net and PLS predictions in Fig. 6. For PLS, the line of best fit for the CV set was guaranteed to be the identity due to the least squares nature of PLS. However, this was not the case for U-Net. Deep learning regressors do not guarantee to find a least squares solution (Igel and Oehmcke 29), especially when training with bulk references and making sample-wise predictions (Engstrøm et al. 19). However, the U-Net’s CV set sYX was very close to its RMSE. Furthermore, bias and scale correction do not change the visual appearance of the chemical maps, nor do they change the values of the ratios in Table 3. For these reasons, no bias or scale corrections were performed.
@@ 4.3 Chemical maps
PLS and the modified U-Net had approximately the same performance when evaluated on the mean fat predictions for pork bellies (Fig. 6). However, as evident by the spatial statistics in Table 3, U-Net generated much more meaningful chemical maps than PLS. In addition to generating spatially aware, locally smooth chemical maps, U-Net almost always stayed inside the $0-100\%$ range. This was unlike PLS, where most predictions lay outside this range.
@@ 4.4 U-Net extensions
While the goal of this study was to serve as a proof of concept, it was likely possible to further decrease the mean squared error of the U-Net by performing more elaborate studies regarding the values of hyperparameters, including the previously mentioned loss-multipliers, learning rate, optimizer, and all other choices that go into training and evaluating a deep learning model. While such performance optimization was not of interest in this study, it is possible to search for more optimal hyperparameters using, for example, grid-search or a Bayesian approach such as a Tree-Structured Parzen Estimator (TPE) (Bergstra et al. 13) as implemented by [1].
In this study, the fat content of pork bellies was analyzed. The dataset includes references for additional parameters (Albano-Gaglio et al. 2). Future developments of this work involve generating chemical maps for multiple parameters using a single model, e.g., a U-Net. A natural extension of the presented loss function would be suitable in this case. First, each parameter should be scaled by its standard deviation to ensure equal contribution to the total loss. Afterward, it is simple to compute an MSE term, Eq. (3), and an SL term, Eq. (5), for each parameter. Interestingly, optimizing the model to regress multiple parameters yields additional possibilities related to the out-of-bounds loss, Eq. (4), that aims to constrain the pixel-wise predictions of each parameter to values between 0 and 100 percent. In addition to constraining each parameter, the sum of parameters should also be constrained in this range. If all parameters can be measured, they should sum to 100 percent. Thus, it is hypothesized that the generalization of the presented work to multiple parameters is not only feasible but perhaps even beneficial compared to the single-parameter case.
Another possible direction for extending this work can be to maintain U-Net’s chemical map generation capability while also having it learn to perform segmentation between foreground (pork belly) and background, a task for which it was initially designed (Ronneberger et al. 41).
@@ 5 Conclusion
This study has proposed a novel end-to-end deep learning approach based on U-Net for chemical map generation using hyperspectral images of pork bellies with associated mean fat reference values as a case study. By training a modified U-Net with a robust optimization scheme and a multi-faceted customized loss function, using only a mean reference value, the model has learned to generate chemical maps with a high degree of spatial correlation and to make predictions strictly inside the 0-100% range of fat values. These findings contrast the current PLS-based pixel-wise approach to chemical map generation that will generate chemical maps with a low degree of spatial correlation and predictions ranging from -2000-2000%, significantly exceeding the realm of the physically possible. Simultaneously, the U-Net-based approach achieves a lower RMSE than PLS when averaging the chemical maps for comparison with the mean reference values. Thus, this study indicates that the U-Net-based approach enables the generation of chemical maps with meaningful spatial features, emphasizing that the analysis of hyperspectral images can benefit from joint consideration of the spectral and spatial features.
@@ Appendix B Loss evolution for U-Net
Appendix C Metrics
Tables 4 and 5 show $Q^{2}$ for the test set and $R^{2}$, RMSE, bias and scale for the CV set and test set. Note that the CV set is everything except the test set and would typically be called the calibration set in chemometric lingo. The bias and scale are the solution to an ordinary least squares regression from the predicted values, $\hat{\mathbf{y}}$ to the reference values, $\mathbf{y}$. They are the values used to compute the lines of best fit in Fig. 6.
Metric
Model $R^{2}$ RMSE bias scale
U-Net 0.85 5.30 -2.90 1.07
PLS 0.91 4.05 0.00 1.00
Table 4: Common chemometric metrics for the models evaluated on the entire CV set (this is everything except the test set and would be called the calibration set in chemometric lingo).
Metric
Model $Q^{2}$ $R^{2}$ RMSE bias scale
U-Net 0.84 0.84 5.25 -0.95 1.03
PLS 0.82 0.82 5.64 4.20 0.91
Table 5: Common chemometric metrics for the models evaluated on the test set.
Appendix D Update
This work is an update of the work by Engstrøm et al. [18], with most of the content remaining similar. This update rectifies a mistake in the original work, as detailed in this appendix. As such, the original work exists only for archival purposes, and this update is the preferred version.
In the original work, an apparent difference was observed between evaluating PLS on the pork belly-wise mean spectra and evaluating PLS as the averages of pixel-wise predictions on the images of the pork bellies. These two operations should be equivalent. However, in the original work, preprocessing with SNV and convolution with Savitzky-Golay filters were applied after computing the mean spectra of pork bellies. Conversely, when evaluating PLS on individual pixels to generate chemical maps, preprocessing was applied directly to those pixels. While convolution with Savitzky-Golay filters yields the same result whether applied on individual pixels before averaging or on the mean spectrum after averaging, SNV does not. In particular, the standard deviation used in the SNV transformation is sensitive to whether it is computed on the mean spectrum or on individual pixels that are subsequently averaged to yield a mean spectrum. Thus, the mean spectra obtained by those two orderings of SNV are different. Thus, in the original work, any discrepancy between evaluating PLS on mean spectra and as the average of pixel-wise predictions was solely due to the order in which SNV was applied. While this procedure was described in the original work in Section 2.2.1, the authors believe that this issue should be resolved; hence, the publication of this update, where SNV is always applied to each pixel individually, prior to any computation of the mean spectra.
Below is an enumeration of the changes made to the original work resulting in this updated version.
  1. 1.
    Change of Section 2.2.1 to indicate that SNV is always applied on the pixel-wise level and, as such, PLS evaluated on mean spectra and as averages of pixel-wise predictions are equivalent. The updated PLS model uses 11 components instead of nine.
  2. 2.
    Removal of any mention of discrepancy between PLS predictions on mean spectra and the averages of pixel-wise PLS predictions.
  3. 3.
    Update of all reports of PLS performance to reflect the updated model’s performance.
  4. 4.
    Update of Fig. 6 to reflect the changes in PLS.
  5. 5.
    Update of Fig. 7 to reflect the changes in PLS.
  6. 6.
    Update of Fig. 8 to reflect the changes in PLS.
  7. 7.
    Update of Fig. 9 to reflect the changes in PLS.
  8. 8.
    Update of Table 2 to reflect the changes in PLS.
  9. 9.
    Update of Table 3 to reflect the changes in PLS.
  10. 10.
    Update of Table 4 to reflect the changes in PLS.
  11. 11.
    Update of Table 5 to reflect the changes in PLS.